OpenStudy (ciel21):

can you help me evaluate the expression when x = –11. 2x – 19

OpenStudy (anonymous):

ok you have to plug it

OpenStudy (anonymous):

plug in -11 to x and solve

OpenStudy (anonymous):

i believe its 8

OpenStudy (anonymous):

2(-11)-19=?

OpenStudy (anonymous):

\[2\times(-11) -19 \to -22-19=-41\]

OpenStudy (anonymous):

11 times 2 equals 22 but since its a negative i believe that u change the sign and add

OpenStudy (anonymous):

oh sorry ya it would be -41 not 11 because my negative turned into a subtract

OpenStudy (ciel21):

i got it it was 2 * -11 - 19

OpenStudy (anonymous):

yes

OpenStudy (ciel21):

:D im so smart

OpenStudy (anonymous):

no i am a good teacher ahah ԅ(≖‿≖ԅ)

OpenStudy (ciel21):

ԅ(≖‿≖ԅ) lies

OpenStudy (anonymous):

hahaah

OpenStudy (ciel21):

ԅ(≖‿≖ԅ) hehehehe..he..hehe.heheheheheeheheheheheheheh